In an individual retirement account valued at $53,900, a couple has 500 shares of stock, some in Big Bank Corporation and some in Safe Savings and Loan. If big bank sells for $115 per share and Safe Savings sells for $97 per share, how many shares of each does the couple own?

(I was thinking 115x+97x=500 , but it did not workout)

Answer by nerdybill(7384)

In an individual retirement account valued at $53,900, a couple has 500 shares of stock, some in Big Bank Corporation and some in Safe Savings and Loan. If big bank sells for $115 per share and Safe Savings sells for $97 per share, how many shares of each does the couple own?
.
Let x = shares of Big Bank
then
500-x = shares of Safe Savings
.
115x + 97(500-x) = 53900
115x + 48500 - 97x = 53900
115x - 97x = 5400
18x = 5400
x = 300 shares (Big Bank)
.
Safe Savings:
500-x = 500-300 = 200 shares